Shelly is rowing a boat upstream. The river is flowing at a speed of 3 miles per hour. Sharon starts rowing at a speed of 8 miles per hour, but as she gets tired her speed decreases (at a rate of 1.5 miles per hour, every hour).
Graph the equation that fits the description, where y represents Sharon's speed and x represents the time elapsed since the start of her workout.

Answer by Theo(13342)

x represents the elapsed time in hours.
y represents the speed in miles per hour.
she starts off at 8 miles per hour and her speed decreased by 1.5 miles per hour every hour.

the equation to model this is y = -1.5 * x + 8

since she is traveling against the current, and the current is traveling at a constant rate of 3 miles per hour, then her overall speed is 3 miles per hour less than if she were rowing in a still current (not with nor against).

the equation to model this is y = -1.5 * x + 8 - 3

simplify this to get y = -1.5 * x + 5.

you can graph this equation and then mark her speed at every hour.

she will be traveling at a net speed of 0 miles per hour in 3 and 1/3 hours.

when she can't row anymore, she will be traveling at 3 miles per hour in the opposite direction (the speed of the current).

the graph looks like this.

$$$

the domain of this graph starts at x = 0 and ends at x = 5 and 1/3.

x can't be less than 0 because time can't be negative.
x can't be greater than 5 and 1/3 because she's completely exhausted and can't row anymore, either forward or backward. she therefore can only drift backwards at the speed of the current.

her speed when she starts is a net 5 miles per hour.
8 miles per hour rowing against 3 miles per hour of current.

when her net speed is 0 miles per hour, she is rowing at a speed of 3 miles per hour against a current of 3 miles per hour.
this occurs in 3 and 1/3 hours.

when her net speed is -3 miles per hour, she is rowing at a speed of 0 miles per hour against a current of 3 miles per hour.
this occurs in 5 and 1/3 hours.

the coordinates of the graph are in (x,y) format.
x represents the hours she has been rowing.
y represents her net speed.

for example:

at (2,2), she has been rowing for x = 2 hours and her net speed is y = 2 miles per hour.

per the formula, y = -1/5 * 2 + 5 which results in y = -3 + 5 resulting in y = 2.